On March 31, 2021, Management of quality appliances committed to a plan to sell equipment. The equipment was available for immediate sale, and an active plan to locate a buyer was initiated. The equipment had been purchased on January 1, 2019 for $260000. The equipment had an estimated six-year service life and a residual value of $20000. The equipment was being depreciated using the straight-line method. Quality's fiscal year ends on December 31.

Prepare the journal entry for the 2021 depreciation expense

What is the equipment's book value as of March 31,2021?

By December 31, 2021 the equipment has not been sold, but management expects that it will be sold in 2022 for $150000. For what amount is the equipment reported in the December 31, 2021 balance sheet. Please prepare appropriate journal entries
Answer #1

a) Journal entry

Date account and explanation Debit Credit
2021 Depreciation expense (260000-20000/6)*3/12 10000
Accumulated depreciation-equipment 10000

b) Accumulated depreciation = (260000-20000/6)*2+10000 = 90000

Book value as of March 31,2021 = 260000-90000 = $170000

c) Equipment reported as on December 31,2021 = 260000-(260000-20000/6*3) = 140000

Date account and explanation Debit Credit
Cash 150000
Accumulated depreciation-equipment 140000
Gain on sale of equipment 30000
equipment 260000